# CoTKR: Chain-of-Thought Enhanced Knowledge Rewriting for Complex Knowledge Graph Question Answering

### v1

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, marking a significant milestone Sanh et al. (2022); Brown et al. (2020); Zhang et al. (2022); Azaria et al. (2024); Chen (2024). Despite their superior performance in zero-shot scenarios Wei et al. (2022a); Kojima et al. (2022), they still encounter factual errors, known as “hallucinations” Ji et al. (2023b), especially in knowledge-intensive tasks like question answering (QA) Hu et al. (2023); Tan et al. (2023); Li et al. (2023). This issue arises due to the intrinsic limitations of LLMs, including factual inaccuracies and outdated knowledge. To address this challenge, a substantial of work Ma et al. (2023a); Trivedi et al. (2023); Wu et al. (2023b) retrieves task-relevant knowledge from external sources as context, thereby enhancing the capabilities of LLMs in downstream tasks, known as Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023).
Recent work Wu et al. (2023b); Baek et al. (2023); Sen et al. (2023); Wang et al. (2024) under the RAG paradigm explores the use of Knowledge Graphs (KGs) as an information source to enhance the capabilities of LLMs in Question Answering (QA). Unlike typical QA tasks, a key challenge in KGQA under this paradigm lies in transforming question-related subgraphs into natural language that LLMs can understand while preserving the structural information Ko et al. (2024); Ding et al. (2024); Wu et al. (2023b). This process is referred to as Knowledge Rewriting (KR) in this study. As illustrated in Figure 1, this paper summarizes the commonly used knowledge rewriting methods in existing work. Most prior studies Baek et al. (2023); Sen et al. (2023); Wang et al. (2023a) employ simple linear concatenation method (Triple), which concatenates the subject, relation, and object of each triple to form triple-form text. Additionally, considering that LLMs are pre-trained on text corpora and struggle with structured triple-form text, some efforts Wu et al. (2023b); Bian et al. (2021); Chen et al. (2022) focus on converting triples into natural language through KG-to-Text. Furthermore, given that retrieved subgraphs often contain redundant information irrelevant to the question, other studies Ko et al. (2024); Dernbach et al. (2024) aim to extract question-relevant knowledge from the triples to generate summary pertinent to the question.
@@ 4.1 Chain-of-Thought Enhanced Knowledge Rewriting
xAssume we have the reasoning traces at step $t-1$ as $x_{t-1,r}$ and the summarized knowledge at step $t-1$ as $x_{t-1,k}$. The corresponding knowledge representation, i.e., $X_{t-1}$ is represented as:
@@ 4.2 Training Framework for CoTKR
Fig. 3 illustrates the training framework of CoTKR.
@@ 4.2.2 Preference Alignment from Question Answering Feedback (PAQAF)
Preference Annotation based on Question Answering Feedback. Among the candidate knowledge representations, we select the two, $k_{1}$ and $k_{2}$, with the greatest semantic difference (i.e., the lowest similarity) to facilitate faster convergence during training. Utilizing standard evaluation methods for assessing these representations is suboptimal, as they fail to align with the preferences of QA models. Inspired by the findings in previous workWu et al. (2023b); Ko et al. (2024); Zhang et al. (2023), we posit that better knowledge representations generally lead to better performance on QA. Consequently, we adopt $k_{1}$ and $k_{2}$ as contextual knowledge, prompting the QA model $Q$ to answer the question $q$, generating answers $a_{1}$ and $a_{2}$, respectively. Subsequently, we prompt ChatGPT to assess the quality of $a_{1}$ and $a_{2}$ from the perspectives of accuracy and relevance. This evaluation aims to identify the preferred knowledge representation $k^{+}$ and the dispreferred knowledge representation $k^{-}$. Details of the evaluation prompt are provided in Appendix B.3.
Data Augmentation based on ChatGPT. ChatGPT is able to produce higher-quality knowledge representations, compared with open-source LLMs. Therefore, in order to improve the quality of preferred knowledge representation and enhance the diversity of the training data, we employ ChatGPT to paraphrase $k^{+}$. In addition to the question $q$, the retrieved subgraph $G^{\prime}$, and the preferred knowledge representation $k^{+}$, we also provide the answer entity $e$. This allows ChatGPT to organize relevant knowledge around $e$, ensuring that the rewritten knowledge covers key evidence. We concatenate the question $q$ and the textualized subgraph $G^{\prime}$ using a prompt template as the input $x$, and use the paraphrased knowledge representation $k^{++}$ and $k^{-}$ as the preferred pair. Finally, we construct the preference dataset $P_{N}={(x_{1},k_{1}^{++},k_{1}^{-}),(x_{2},k_{2}^{++},k_{2}^{-}),...,(x_{N},k_{N}^{++},k_{N}^{-})}$. The prompt for knowledge augmentation is in Appendix B.4.
@@ 5.5 Evaluation Metrics
Following previous work on generative KGQA Wu et al. (2023b); Baek et al. (2023); Ko et al. (2024), we adopt Accuracy (Acc) as one of our evaluation metrics. It measures whether the generated answer includes at least one correct answer entity. To provide a more comprehensive evaluation, we also incorporate Recall which measures the proportion of correct entities present in the model’s response, and Exact Match (EM) which evaluates whether the response contains all the answer entities.
@@ 5.9 Effectiveness of Data Augmentation
Table 2: The ablation results of data augmentation on GraphQuestions. We use 2-Hop as the retrieval method.
@@ 5.10 Case Study
In this section, we compare Summary with CoTKR through an example. (Please refer to Appendix C for the full comparison result.) As illustrated in Table 3, Summary overlooks the key knowledge (i.e., Stephen Jay Gould). In contrast, CoTKR highlights the most relevant knowledge in the first rewriting step (i.e., Knowledge 1). Furthermore, after preference alignment, CoTKR+PA can avoid the unnecessary reasoning step (i.e., Reason 2), generating comprehensive and semantic coherent contextual knowledge for QA.
Question: what is the central focus of the radio program episode talk of the nation: memes?
Answer: Stephen Jay Gould
KR Methods KR Results
Summary The central focus of the radio program episode "Talk of the Nation: Memes" is discussing topics related to memes and their impact on society.
[…]
Summarize: The subject of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
Reason: I need to know more about Stephen Jay Gould.
Summarize: Stephen Jay Gould was a person with nationality …
[…]
Summarize: The subject of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
@@ Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. U21A20488 and by the project "Key Laboratory of rich-media Digital Publishing Content Organization and Knowledge Service Open Fund-Research on Knowledge-enhanced Training Techinques of Large Language Model" No. ZD2024-04/01. Jiaoyan Chen is funded by the EPSRC project OntoEm (EP/Y017706/1). We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper. ChatGPT was used to enhance the readability of some of the text and improve the language of this paper, after the content was first added manually. All material was then checked manually before submission.
@@ A.3 Large Language Models
A.4 Retrieval Methods Details
A.5 Implementation Details
Appendix B Prompts
In this section, we provide the prompts (omit demonstrations) used in the experiment.
B.1 Prompts for Knowledge Rewriting Methods
Table 4, Table 5, and Table 6 show the prompts for different knowledge rewriting methods.
@@ B.1 Prompts for Knowledge Rewriting Methods
Table 6: CoTKR Prompt.
B.2 Prompt for Question Answering
Table 7, Table 8, and Table 9 show the prompts for question answering.
@@ B.2 Prompt for Question Answering
Table 7: Prompt for Question Answering with Triple/KG-to-Text/Summary Knowledge.
@@
Table 8: Prompt for Question Answering with CoTKR/CoTKR+PA Knowledge.
@@
Table 9: Prompt for Question Answering without Context.
B.3 Prompt for Preference Annotation
Table 10 shows the preference annotation prompt.
@@ B.3 Prompt for Preference Annotation
B.4 Prompt for Data Augmentation
Table 11 shows the data augmentation prompt.
@@ B.4 Prompt for Data Augmentation
Appendix C Case Study
In this section, we compare different knowledge rewriting strategies through an example. As illustrated in Table 12, the knowledge generated by both Triple and KG-to-Text contains excessive redundant information. This redundancy complicates the process for the QA model, making it challenging to extract relevant knowledge. Summary omits crucial information (i.e., Stephen Jay Gould), rendering the QA model almost unable to provide the correct answer. In contrast, CoTKR and CoTKR+PA emphasize the most pertinent knowledge in the first rewriting step, thereby enabling the QA model to provide a concise and accurate answer. Furthermore, after preference alignment, our knowledge rewriter avoids unnecessary inferential steps (i.e., Reason 2), generating comprehensive and semantic coherent contextual knowledge for QA.
Question: what is the central focus of the radio program episode talk of the nation: memes?
Answer: Stephen Jay Gould
@@ Appendix C Case Study
Triple (Talk of the Nation: Memes, radio.radio_program_episode.program, Talk of the Nation) (Talk of the Nation: Memes, radio.radio_program_episode.subjects, Stephen Jay Gould) (Talk of the Nation, broadcast.content.artist, Neal Conan) (Talk of the Nation, broadcast.content.genre, News)… The central focus of the radio program episode Talk of the Nation: Memes is the discussion of memes, which are cultural ideas, symbols, or practices that spread from person to person within a culture.
@@
Summarize: The subject of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
Reason: I need to know more about Stephen Jay Gould.
Summarize: Stephen Jay Gould was a person with nationality…
The central focus of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
@@
Summarize: The subject of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
The central focus of the radio program episode "Talk of the Nation: Memes" is Stephen Jay Gould.
@@
Table 12: An example of knowledge rewriting results for different methods. We use Llama-2 and ChatGPT as the Knowledge Rewriting (KR) backbone and the QA model, respectively.
Appendix D Additional Experimental Results

### v2

@@ 1 Introduction
Large Language Models (LLMs) have achieved remarkable performance across various natural language processing tasks, marking a significant milestone Sanh et al. (2022); Brown et al. (2020); Zhang et al. (2022); Azaria et al. (2024); Chen (2024). Despite their superior performance in zero-shot scenarios Wei et al. (2022a); Kojima et al. (2022), they still encounter factual errors, known as “hallucinations” Ji et al. (2023b), especially in knowledge-intensive tasks like question answering (QA) Hu et al. (2023); Tan et al. (2023); Li et al. (2023); He et al. (2023). This issue arises due to the intrinsic limitations of LLMs, including factual inaccuracies and outdated knowledge Pan et al. (2023). To address this challenge, a substantial of work Ma et al. (2023a); Trivedi et al. (2023); Wu et al. (2023b) retrieves task-relevant knowledge from external sources as context, thereby enhancing the capabilities of LLMs in downstream tasks, known as Retrieval-Augmented Generation (RAG) Lewis et al. (2020); Gao et al. (2023); Huang et al. (2023).
@@
Recent work Wu et al. (2023b); Baek et al. (2023); Sen et al. (2023); Wang et al. (2024) under the RAG paradigm explores the use of Knowledge Graphs (KGs) Pan et al. (2017b, a) as an information source to enhance the capabilities of LLMs in Question Answering (QA). Unlike typical QA tasks, a key challenge in KGQA under this paradigm lies in transforming question-related subgraphs into natural language that LLMs can understand while preserving the structural information Ko et al. (2024); Ding et al. (2024); Wu et al. (2023b). This process is referred to as Knowledge Rewriting (KR) in this study. As illustrated in Figure 1, this paper summarizes the commonly used knowledge rewriting methods in existing work. Most prior studies Baek et al. (2023); Sen et al. (2023); Wang et al. (2023a) employ simple linear concatenation method (Triple), which concatenates the subject, relation, and object of each triple to form triple-form text. Additionally, considering that LLMs are pre-trained on text corpora and struggle with structured triple-form text, some efforts Wu et al. (2023b); Bian et al. (2021); Chen et al. (2022) focus on converting triples into natural language through KG-to-Text. Furthermore, given that retrieved subgraphs often contain redundant information irrelevant to the question, other studies Ko et al. (2024); Dernbach et al. (2024) aim to extract question-relevant knowledge from the triples to generate summary pertinent to the question.
@@ 4.1 Chain-of-Thought Enhanced Knowledge Rewriting
Assume we have the reasoning traces at step $t-1$ as $x_{t-1,r}$ and the summarized knowledge at step $t-1$ as $x_{t-1,k}$. The corresponding knowledge representation, i.e., $X_{t-1}$ is represented as:
@@ 4.2 Training Framework for CoTKR
Figure 3 illustrates the training framework of CoTKR.
@@ 4.2.2 Preference Alignment from Question Answering Feedback (PAQAF)
Preference Annotation based on Question Answering Feedback. Among the candidate knowledge representations, we select the two, $k_{1}$ and $k_{2}$, with the greatest semantic difference (i.e., the lowest similarity) to facilitate faster convergence during training. Utilizing standard evaluation methods for assessing these representations is suboptimal, as they fail to align with the preferences of QA models. Inspired by the findings in previous workWu et al. (2023b); Ko et al. (2024); Zhang et al. (2023), we posit that better knowledge representations generally lead to better performance on QA. Consequently, we adopt $k_{1}$ and $k_{2}$ as contextual knowledge, prompting the QA model $Q$ to answer the question $q$, generating answers $a_{1}$ and $a_{2}$, respectively. Subsequently, we prompt ChatGPT to assess the quality of $a_{1}$ and $a_{2}$ from the perspectives of accuracy and relevance. This evaluation aims to identify the preferred knowledge representation $k^{+}$ and the dispreferred knowledge representation $k^{-}$. Details of the evaluation prompt are provided in Appendix 8.
Data Augmentation based on ChatGPT. ChatGPT is able to produce higher-quality knowledge representations, compared with open-source LLMs. Therefore, in order to improve the quality of preferred knowledge representation and enhance the diversity of the training data, we employ ChatGPT to paraphrase $k^{+}$. In addition to the question $q$, the retrieved subgraph $G^{\prime}$, and the preferred knowledge representation $k^{+}$, we also provide the answer entity $e$. This allows ChatGPT to organize relevant knowledge around $e$, ensuring that the rewritten knowledge covers key evidence. We concatenate the question $q$ and the textualized subgraph $G^{\prime}$ using a prompt template as the input $x$, and use the paraphrased knowledge representation $k^{++}$ and $k^{-}$ as the preferred pair. Finally, we construct the preference dataset $P_{N}={(x_{1},k_{1}^{++},k_{1}^{-}),(x_{2},k_{2}^{++},k_{2}^{-}),...,(x_{N},k_{N}^{++},k_{N}^{-})}$. The prompt for knowledge augmentation is in Appendix A.6.
@@ 5.5 Evaluation Metrics
Following previous work on generative KGQA Wu et al. (2023b); Baek et al. (2023); Ko et al. (2024), we adopt Accuracy (Acc) as one of our evaluation metrics. It measures whether the model’s response includes at least one answer entity. For a dataset comprising $N$ questions, Acc is calculated as follows:
$Acc=\frac{\sum_{i=1}^{N}Acc_{i}}{N}$ (7)
$Acc_{i}=\begin{cases}1,&\text{if at least one answer entity}\\ &\text{appears in the response}\\ 0,&\text{if no answer entity appears in}\\ &\text{the response}\end{cases}$ (8)
In order to comprehensively assess the performance of KGQA, we employ Recall to evaluate the proportion of correct answer entities present in the model’s response. For a dataset containing $N$ questions, Recall is calculated as follows:
$Recall=\frac{\sum_{i=1}^{N}Recall_{i}}{N}$ (9)
$Recall_{i}=\frac{N_{appear}}{N_{total}}$ (10)
where $N_{appear}$ refers to the number of answer entities contained in the model’s responses, and $N_{total}$ refers to the total number of the answer entities. Additionally, we consider utilizing Exact Match (EM) as an evaluation metric. Given that the responses generated by LLMs consist of multiple paragraphs, while the corresponding answers are entities, we adjust the traditional EM metric. Our modified EM metric assesses whether all answer entities are included in the model’s responses. For a dataset consisting of $N$ questions, EM is calculated as follows:
$EM=\frac{\sum_{i=1}^{N}EM_{i}}{N}$ (11)
$EM_{i}=\begin{cases}1,&\text{if all answer entities}\\ &\text{appear in the response}\\ 0,&\text{other cases}\end{cases}$ (12)
@@ 5.9 Effectiveness of Data Augmentation
Table 2: Ablation study results for data augmentation on GraphQuestions, employing the 2-Hop retrieval method.
Question: what is the unit of area that the measurement system that have an electric field strength units of volt per metre have?
Answer: Square kilometer, Square meter
KR Methods KR Results
Summary The International System of Units includes various measurement systems for different physical quantities such as length, area, volume, conductance, resistance, current, force, and pressure. The measurement system with an electric field strength unit of Volt per meter does not directly relate to a specific area unit within the International System of Units.
[…]
Summarize: The International System of Units includes various units for different physical quantities such as length, area, volume, etc.
Reason: I need to find the specific unit of area in the International System of Units that corresponds to the electric field strength unit of Volt per meter.
Summarize: The International System of Units includes units like Square kilometer, Square meter, etc. for measuring area.
[…]
Knowledge: The International System of Units contains the unit for electric field strength in Volt per meter.
Reason: I need to identify the unit of area within the International System of Units.
Knowledge: The International System of Units includes the units Square kilometer and Square meter as area units.
@@ 5.10 Case Study
In this section, we compare Summary with CoTKR through an example. (Please refer to Appendix B for the full comparison result.) As illustrated in Table 3, Summary struggles to extract useful information when faced with an abundance of triples. In contrast, CoTKR, leveraging CoT reasoning, effectively emphasizes the key evidence (i.e., Square meter) in the second rewriting step. Furthermore, after preference alignment, CoTKR+PA is capable of generating more natural reasoning steps, significantly enhancing its applicability to KGQA.
@@ Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. U21A20488, by the project "Key Laboratory of rich-media Digital Publishing Content Organization and Knowledge Service Open Fund-Research on Knowledge-enhanced Training Techinques of Large Language Model" No. ZD2024-04/01, and by the EPSRC project OntoEm (EP/Y017706/1). We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper. ChatGPT was used to enhance the readability of some of the text and improve the language of this paper, after the content was first added manually. All material was then checked manually before submission.
@@ A.3 Large Language Models
Table 4: Prompts for Knowledge Rewriting Methods.
@@
Table 5: Prompts for Question Answering.
@@
A.4 Retrieval Methods Details
@@
A.5 Implementation Details
Question: what is the unit of area that the measurement system that have an electric field strength units of volt per metre have?
Answer: Square kilometer, Square meter
@@ A.5 Implementation Details
Triple (International System of Units, electric_field_strength_units, Volt per metre) (International System of Units, area_units, Square meter) (International System of Units, area_units, Square kilometer) (Volt per metre, strength_in_volts_per_meter, 1.0) … Square meter.
@@
Summarize: The International System of Units includes various units for different physical quantities such as length, area, volume, etc.
Reason: I need to find the specific unit of area in the International System of Units that corresponds to the electric field strength unit of Volt per meter.
Summarize: The International System of Units includes units like Square kilometer, Square meter, etc. for measuring area.
The unit of area in the measurement system that has an electric field strength unit of Volt per meter is Square meter (m²).
@@
Knowledge: The International System of Units contains the unit for electric field strength in Volt per meter.
Reason: I need to identify the unit of area within the International System of Units.
Knowledge: The International System of Units includes the units Square kilometer and Square meter as area units.
The unit of area within the International System of Units that the measurement system having an electric field strength unit of volt per metre has is either Square kilometer or Square meter.
@@
Table 8: An example of knowledge rewriting results for different methods. We use Llama-3 and Mistral as the Knowledge Rewriting (KR) backbone and the QA model, respectively.
All the prompts involved in this experiment are as follows. Table 4 shows the prompts for different knowledge rewriting methods. Table 5 shows the prompts for question answering. Table 6 shows the preference annotation prompt.
A.6 Prompt for Data Augmentation
Table 7 shows the data augmentation prompt.
Appendix B Case Study
In this section, we compare different knowledge rewriting strategies through an example. As illustrated in Table 8, the knowledge generated by both Triple and KG-to-Text contains excessive redundant information. This redundancy complicates the process for the QA model, making it challenging to extract relevant knowledge. Summary struggles to extract useful information when faced with an abundance of triples. In contrast, CoTKR and CoTKR+PA summarize the most pertinent knowledge in the rewriting step, thereby enabling the QA model to provide a concise and accurate answer. Furthermore, after preference alignment, our knowledge rewriter is capable of generating more natural reasoning steps, significantly enhancing its applicability to KGQA.
Appendix C Additional Experimental Results
C.1 Experiments on ComplexWebQuestions
To evaluate the robustness of CoTKR, we conduct our experiments on ComplexWebQuestionsTalmor and Berant (2018). We utilize ChatGPT as the knowledge rewriter, Mistral as the question-answering model, and 2-Hop as the retrieval method. The experimental results, presented in Table 9, demonstrate the effectiveness of CoTKR.
Methods Acc Recall EM
No Knowledge 36.82 31.30 27.62
Triple 39.29 33.94 30.86
KG-to-Text 35.96 31.53 28.96
Summary 38.53 33.87 30.89
CoTKR 40.70 35.74 32.72
Table 9: Experiments on ComplexWebQuestions use ChatGPT as the knowledge rewriter, Mistral as the QA model, and 2-Hop for retrieval.
C.2 Knowledge Rewriter with GPT-4
To assess the applicability of CoTKR to GPT-4, we further conduct the experiments with GPT-4 as the knowledge rewriter, Mistral as the question-answering model, and 2 hop as the retrieval method on 1,000 test questions from GrailQA. The detailed results are presented in Table 10. The findings show that CoTKR outperforms other approaches, with CoTKR utilizing GPT-4 achieving the highest performance. This suggests that employing a more advanced LLM backbone, such as GPT-4, leads to superior outcomes.
Methods Acc Recall EM
No Knowledge 29.10 21.87 18.80
Triple 54.30 47.15 42.40
KG-to-Text 53.20 45.42 40.60
Summary 56.00 48.63 43.60
CoTKR 57.50 52.16 48.20
Table 10: Experiments on 1,000 GrailQA test questions use GPT-4 for rewriting, Mistral for QA, and 2-Hop as the retrieval method.
C.3 Time Analysis
We conduct experiments to analyze the average time cost of the knowledge rewriting methods discussed in this paper. We adopt Llama-3 as the knowledge rewriter, Mistral as the question-answering model, and 2-Hop as the retrieval method. The experiments are conducted on GraphQuestions, utilizing one A100-SXM4-40GB GPU. The average runtime for each question by different methods is shown in Table 11 (unit: seconds). The average runtime of each question for all methods is within an acceptable range (i.e., less than 1.5 seconds). Although our method is the most time-consuming, it exhibits a clear advantage in performance.
Process No Knowledge Triple KG-to-Text Summary CoTKR
Rewrite 0 0 1.0828 0.4615 0.9835
Answer 0.3787 0.5998 0.2701 0.2502 0.4414
Rewrite+Answer 0.3787 0.5998 1.3529 0.7117 1.4249
Table 11: Time analysis on GraphQuestions (seconds) using Llama-3 as the knowledge rewriter, Mistral for question answering, and 2-Hop for retrieval.
C.4 GPT-4-score as Evaluation Metrics
Given the powerful natural language understanding and generation capabilities of closed-source large models, many existing works employ ChatGPT as an evaluator to provide high-quality evaluation results Sottana et al. (2023); Liu et al. (2023); Min et al. (2023). In our approach, we use GPT-4 as the evaluator to assess whether all answer entities are present in the responses. We refer to this evaluation metric as GPT-4-score. Compared to EM, this metric is more flexible, as LLMs are capable of recognizing synonyms of answer entities. We use it to evaluate the first 300 questions from GrailQA using Llama-3 as knowledge rewriter, ChatGPT as question-answering model, and 2 hop as retrieval method. We provide the prompt for GPT-4-score in Table 12.
GPT-4-score Prompt
[Instruction]
Your task is to evaluate the quality of the response to the question. You should consider whether all the answer entities appear in the response.
Question: {question} Answer: {answer} Response: {response} Begin your evaluation by comparing the response and the answer and provide a short explanation. Then output only the single number: "1" if all the answer entities appear in the response, and "0" if not. At the end, repeat just the number again by itself on a new line.
Table 12: GPT-4-score Prompt.
The experimental results are shown in Table 13. The results indicate that our implemented GPT-4-score is effective and consistent with the outcomes reflected by other evaluation metrics. Furthermore, this also demonstrates that CoTKR possesses significant advantages compared to other knowledge rewriting methods.
Methods GPT-4-score
No Knowledge 0.2767
Triple 0.4700
KG-to-Text 0.4749
Summary 0.4800
CoTKR 0.5167
CoTKR+PA 0.5567
Table 13: GPT-4-score on GraphQuestions. We use Llama-3 as knowledge rewriter, ChatGPT as question-answering model, and 2 hop as retrieval method.
C.5 Qualitative Analysis on Data Augmentation
To clearly demonstrate the importance of data augmentation, we perform a qualitative analysis. The results reveal that data augmentation enhances performance in three key areas: reducing redundant reasoning steps, supplementing critical information, and generating more concise summaries. These improvements are exemplified by three representative cases, as shown in Table 14.
Effect Question $k^{+}$ $k^{++}$
Avoiding Redundant Reasoning Step Katy börner is the curator for what exhibition? Reason: I need to know which exhibition Katy Börner curated.
Summarize: Katy Börner curated the exhibition "Places & Spaces".
Reason: I need to know the subjects of the exhibition "Places & Spaces" curated by Katy Börner.
Summarize: The subjects of the exhibition "Places & Spaces" curated by Katy Börner include Cartography, Information visualization, Information design, and Scientometrics.
Reason: I need to know which exhibition Katy Börner curated.
Summarize: Katy Börner curated the exhibition "Places & Spaces".
Key Points Supplementation Person of interest deals with how many tv subjects? Reason: I need to know how many TV subjects are associated with "Person of Interest".
Summarize: "Person of Interest" is associated with multiple TV subjects such as Espionage, Mass surveillance, Hacker, and Vigilante.
Reason: I need to know how many TV subjects are associated with "Person of Interest".
Summarize: "Person of Interest" deals with a total of 5 TV subjects including Espionage, Mass surveillance, Hacker, Surveillance, and Vigilante.
More Concise Summarization Who are the owners of gree? Reason: I need to know who owns GREE, Inc.
Summarize: GREE, Inc. is owned by GREE, Inc. itself, according to the triple (GREE, internet.website.owner, GREE, Inc.).
Reason: I need to know who owns GREE, Inc.
Summarize: GREE, Inc. is owned by GREE, Inc. itself.
Table 14: Representative cases for improvements through data augmentation.
C.6 Detailed Experimental Results